OpenStudy (anonymous):

How do you factor x^2+4?

OpenStudy (anonymous):

(x+2)(x+2)

OpenStudy (anonymous):

actually that is not factorable

OpenStudy (anonymous):

yes i agree with adrian

OpenStudy (anonymous):

They have no factors in common.

OpenStudy (anonymous):

was it x^2-4, it'd be (x+2)(x-2)

OpenStudy (anonymous):

ah that's what I thought..so I'm trying to find the domain of an equation and I'm here x^2+4=0 now what?

myininaya (myininaya):

x^2+4=(x-2i)(x+2i)

myininaya (myininaya):

\[a^2+b^2=(\sqrt{a^2}-\sqrt{-b^2})(\sqrt{a^2}+\sqrt{-b^2})=(a-i b)(a+i b)\]

myininaya (myininaya):

\[(a-ib)(a+ib)=a^2+abi-abi-i^2b^2=a^2-(-1)b^2=a^2+b^2\]